# Positive traces on certain $\operatorname{SL}(2)$ Coulomb branches

### v1

@@ 1.1 Preliminaries
Throughout the chapter, we fix a positive even integer $n$ with $m=\frac{n}{2}$ and denote $\varepsilon=\exp(\frac{2\pi i}{n})$. For a finite subgroup $G$ of the special linear group $\operatorname{SL}_{2}(\mathbb{C})$, the Kleinian singularity $\mathbb{C}^{2}/G$ has as its coordinate ring the commutative graded algebra $A=\mathbb{C}[u,v]^{G}$ of $G$-invariant polynomials in two variables, with the action of $G$ on $A$ defined by
@@
$g=\begin{bmatrix}\varepsilon&0\\ 0&\varepsilon^{-1}\end{bmatrix},\quad\quad h=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$
@@
All elements of the algebra $A$ are the sum of homogeneous polynomials, so $A$ is graded by the degree of homogeneous polynomials. We consider filtered deformations $\mathcal{A}$ of $A$, that is, filtered associative algebras $\mathcal{A}$ with associated graded algebra $\operatorname{gr}\mathcal{A}$ equal to $A$. In particular, we make use of the Crawley–Boevey—Holland CBH (98) filtered deformations. Let $\mathbb{C}[x,y]\#G$ be the skew group algebra constructed from the vector space $\mathbb{C}[x,y]\otimes\mathbb{C}[G]$ with multiplication given by
$(P\otimes g)(G\otimes h)=Pg(Q)\otimes gh,$
@@
For the idempotent $e=\frac{1}{|G|}\sum_{g\in G}g$, let $\mathcal{O}_{c}=eH_{c}e$, which is an algebra with unit $e$. Losev (Los, 22, Theorem 3.4) proved that for any finite subgroup $G$ of $\operatorname{SL}_{2}(\mathbb{C})$, any filtered deformation of $\mathbb{C}[x,y]^{G}$ is isomorphic to some algebra $\mathcal{O}_{c}$ with $c\in Z(\mathbb{C}[g])$.
A trace on $\mathcal{A}$ is a linear function $T\colon\mathcal{A}\rightarrow\mathbb{C}$ satisfying $T(ab)=T(ba)$ for all $a,b\in\mathcal{A}$. We only consider traces on noncommutative algebras, so by (CBH, 98, Theorem 0.4(1)), we may assume the coefficient of $c$ on the group identity is nonzero. Also, by (CBH, 98, Theorem 0.4(2)) that for $c$ outside a finite set of hyperplanes, any trace on $\mathcal{O}_{c}$ is a restriction of a trace on $H_{c}$. We consider only such $c$, which we call generic. For a classification of traces on $\mathcal{O}_{c}$, it therefore suffices to characterize traces on $H_{c}$.
@@
  • $gx=\varepsilon xg$;
  • $gy=\varepsilon^{-1}yg$;
  • $hx=yh$;
  • $hy=-xh$;
  • $xy-yx=c=\sum_{i=0}^{n-1}c_{i}g^{i}$, with $c_{0}\neq 0$;
  • $c_{i}=c_{-i}$ for all $i$.
Here the equality $c_{i}=c_{-i}$ holds because $c$ is in $Z(\mathbb{C}[\mathbb{D}_{n}])$ and the conjugacy classes of $\mathbb{D}_{n}$ are of the form $\{g^{i},g^{-i}\}$, $\{g^{2i}h:i\in\mathbb{Z}\}$, and $\{g^{2i+1}h:i\in\mathbb{Z}\}$.
We close this section with a definition of positive traces. For an algebra $B$ and an antilinear automorphism $\rho\colon B\to B$ such that $\rho^{2}=\operatorname{id}$, a trace $T$ on $B$ is said to be positive if $T(a\rho(a))>0$ for all nonzero $a$ in $B$. The antilinear automorphism we study is defined on $H_{c}$ by $\rho(x)=y$, $\rho(y)=-x$, $\rho(g)=g$, $\rho(h)=h$. We return to positive traces in Section 1.5.
1.2 Algebra Identities and Relations
In this section, we introduce several important algebra elements and describe relations between them.
Recall that $c=\sum_{i=0}^{n-1}c_{i}g^{i}$, with $g$ a generator of $\mathbb{D}_{n}$ such that $g^{n}$ is the group identity, and $\varepsilon=\exp(\frac{2\pi i}{n})$. Let
@@ 1.2 Algebra Identities and Relations
$hk$ $=\frac{1}{c_{0}}\left(hxy+\sum\limits_{i=1}^{n-1}\frac{hc_{i}g^{i}}{\varepsilon^{i}-1}\right)-\frac{h}{2}$
$=\frac{1}{c_{0}}\left(-yxh+\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{-i}h}{\varepsilon^{i}-1}\right)-\frac{h}{2}$
$=\frac{1}{c_{0}}\left(c-xy+\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}}{\varepsilon^{-i}-1}\right)h-\frac{h}{2}$
$=\frac{1}{c_{0}}\left(-xy+c_{0}+\sum\limits_{i=1}^{n-1}\left(c_{i}g^{i}-\frac{\varepsilon^{i}c_{i}g^{i}}{\varepsilon^{i}-1}\right)\right)h-\frac{h}{2}$
$=\frac{1}{c_{0}}\left(-xy+c_{0}-\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}}{\varepsilon^{i}-1}\right)h-\frac{h}{2}$
$=-\frac{1}{c_{0}}\left(xy+\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}}{\varepsilon^{i}-1}\right)h+\frac{h}{2}$
$=-kh.$
@@
$[k,x]$ $=\frac{1}{c_{0}}\left(xyx-x^{2}y+\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}x-xc_{i}g^{i}}{\varepsilon^{i}-1}\right)$
$=\frac{1}{c_{0}}\left(-xc+\sum\limits_{i=1}^{n-1}\frac{\varepsilon^{i}xc_{i}g^{i}-xc_{i}g^{i}}{\varepsilon^{i}-1}\right)$
$=\frac{1}{c_{0}}\left(-xc+x\sum\limits_{i=1}^{n-1}c_{i}g^{i}\right)$
$=-x.$
It follows immediately that $kx=x(k-1)$, and the commutation relations with $y$ follow by conjugating the commutation relations with $x$ by $h$. ∎
@@
Clearly $g^{i}$ and $e_{q}$ commute. We have
@@
$he_{q}=\frac{1}{n}\sum\limits_{i=0}^{n-1}\varepsilon^{iq}hg^{i}=\frac{1}{n}\sum\limits_{i=0}^{n-1}\varepsilon^{iq}g^{-i}h=\frac{1}{n}\sum\limits_{i=0}^{n-1}\varepsilon^{-iq}g^{i}h=e_{-q}h,$
@@
$e_{q}x=\frac{1}{n}\sum\limits_{i=0}^{n-1}\varepsilon^{iq}g^{i}x=\frac{1}{n}\sum\limits_{i=0}^{n-1}\varepsilon^{iq+i}xg^{i}=\frac{1}{n}x\sum\limits_{i=0}^{n-1}\varepsilon^{i(q+1)}g^{i}=xe_{q+1},$
@@
$\alpha_{q}$ $=\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{-iq}}{\varepsilon^{i}-1}=\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{-i}\varepsilon^{-iq}}{\varepsilon^{i}-1}=\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{iq}}{\varepsilon^{-i}-1}$
$=\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{i(q+1)}}{1-\varepsilon^{i}}=-\alpha_{-q-1}.$
@@
$\alpha_{p}e_{q}$ $=\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{-ip}e_{q}}{\varepsilon^{i}-1}$
$=\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{i(q-p)}\varepsilon^{-iq}e_{q}}{\varepsilon^{i}-1}$
$=\frac{1}{c_{0}}\left(\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{i(q-p)}g^{i}}{\varepsilon^{i}-1}\right)e_{q},$
@@
$k=\frac{1}{c_{0}}\left(xy-\sum\limits_{q=0}^{n-1}\beta_{q}e_{q}\right),\quad\textrm{or}\quad xy=c_{0}k+\sum\limits_{q=0}^{n-1}\beta_{q}e_{q},$
@@
$x^{a}y^{a}$ $=\prod_{i=0}^{a-1}\left(c_{0}(k+i)+\sum\limits_{q=0}^{n-1}\beta_{q+i}e_{q}\right),$
$y^{a}x^{a}$ $=(-1)^{a}\prod\limits_{i=0}^{a-1}\left(c_{0}(-k+i)+\sum\limits_{q=0}^{n-1}\beta_{q+i}e_{-q}\right).$
Also, the elements $x^{a}y^{a}e_{q}$ and $y^{a}x^{a}e_{q}$ can be written as polynomial expressions in $k$ multiplied by $e_{q}$ as follows:
@@
We prove by induction the polynomial expression for $x^{a}y^{a}$ by induction, with the base case of $a=1$ true by definition. Assume that $x^{a}y^{a}e_{q}=e_{q}\prod_{i=0}^{a-1}(c_{0}(k+i)+\sum_{q=0}^{n-1}\beta_{q+i}e_{q})$. Then
$x^{a+1}y^{a+1}$ $=x^{a}(xy)y^{a}$
$=x^{a}\left(c_{0}k+\sum\limits_{q=0}^{n-1}\beta_{q}e_{q}\right)y^{a}$
$=x^{a}y^{a}\left(c_{0}(k+a)+\sum\limits_{q=0}^{n-1}\beta_{q}e_{q-a}\right)$
$=x^{a}y^{a}\left(c_{0}(k+a)+\sum\limits_{q=0}^{n-1}\beta_{q+a}e_{q}\right)$
$=\left(\prod\limits_{i=0}^{a-1}\left(c_{0}(k+i)+\sum\limits_{q=0}^{n-1}\beta_{q+i}e_{q}\right)\right)\left(c_{0}(k+a)+\sum\limits_{q=0}^{n-1}\beta_{q+a}e_{q}\right)$
$=\prod\limits_{i=0}^{a}\left(c_{0}(k+i)+\sum\limits_{q=0}^{n-1}\beta_{q+i}e_{q}\right),$
as needed. The analogous claim for $y^{a}x^{a}$ follows from conjugation by $h$. The expressions for $x^{a}y^{a}e_{q}$ and $y^{a}x^{a}e_{q}$ follow using the statement that $e_{p}e_{q}=\delta_{p,q}e_{q}$ from Lemma 1.2.2. ∎
@@
By Lemma 1.2.3, we have
$xye_{q}$ $=\left(c_{0}k+\frac{c_{0}}{2}-\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}}{\varepsilon^{i}-1}\right)e_{q}$
$=\left(c_{0}k+\frac{c_{0}}{2}-c_{0}\alpha_{q}\right)e_{q}.$
Also, by definition, $xye_{q}=\left(c_{0}k+\beta_{q}\right)e_{q}$. Thus
$\left(c_{0}k+\beta_{q}\right)e_{q}$ $=\left(c_{0}k+\frac{c_{0}}{2}-c_{0}\alpha_{q}\right)e_{q}$
$\beta_{q}e_{q}$ $=\left(\frac{c_{0}}{2}-c_{0}\alpha_{q}\right)e_{q},$
@@
1.3 Characterization of Traces
In this section, we study traces using basis elements of the form $x^{i}R(k)e_{q}$, $y^{i}R(k)e_{q}$, $x^{i}R(k)e_{q}h$, and $y^{i}R(k)e_{q}h$, with $R(k)\in\mathbb{C}[k]$. We begin with a lemma about the values $T(e_{q}h)$.
Lemma 1.3.1.
If $m\nmid q$, then the value $T(e_{q}h)$ is equal to 0.

### v2

@@ 2 Preliminaries
Throughout Sections 26, we fix a positive even integer $n$ with $m=\frac{n}{2}$ and denote $\varepsilon=\exp\bigl(\frac{2\pi{\rm i}}{n}\bigr)$. For a finite subgroup $G$ of the special linear group $\operatorname{SL}_{2}(\mathbb{C})$, the Kleinian singularity $\mathbb{C}^{2}/G$ has as its coordinate ring the commutative graded algebra $A=\mathbb{C}[u,v]^{G}$ of $G$-invariant polynomials in two variables, with the action of $G$ on $A$ defined by
@@
$g=\begin{bmatrix}\varepsilon&0\\ 0&\varepsilon^{-1}\end{bmatrix},\qquad h=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$
@@
All elements of the algebra $A$ are sums of homogeneous polynomials, so $A$ is graded by the degree of homogeneous polynomials. We consider filtered deformations $\mathcal{A}$ of $A$, that is, filtered associative algebras $\mathcal{A}$ with associated graded algebra $\operatorname{gr}\mathcal{A}$ equal to $A$. In particular, we make use of the Crawley–Boevey–Holland [6] filtered deformations. Let $\mathbb{C}[x,y]\#G$ be the skew group algebra constructed from the vector space $\mathbb{C}[x,y]\otimes\mathbb{C}[G]$ with multiplication given by
$(P\otimes g)(Q\otimes h)=Pg(Q)\otimes gh,$
@@
For the idempotent $e=\frac{1}{|G|}\sum_{g\in G}g$, let $\mathcal{O}_{c}=eH_{c}e$, which is an algebra with unit $e$. Losev [19, Theorem 3.4] proved that for any finite subgroup $G$ of $\operatorname{SL}_{2}(\mathbb{C})$, any filtered deformation of $\mathbb{C}[x,y]^{G}$ is isomorphic to some algebra $\mathcal{O}_{c}$ with $c\in Z(\mathbb{C}[G])$.
Definition 2.1.
A trace on $\mathcal{A}$ is a linear function $T\colon\mathcal{A}\rightarrow\mathbb{C}$ satisfying $T(ab)=T(ba)$ for all $a,b\in\mathcal{A}$}.
We only consider traces on noncommutative algebras, so by [6, Theorem 0.4 (1)], we may assume the coefficient of $c$ on the group identity is nonzero. Also, by [6, Theorem 0.4 (2)] that for $c$ outside a finite set of hyperplanes, the algebra $\mathcal{O}_{c}$ is Morita equivalent to $H_{c}$. We consider only such $c$, which we call generic. We claim that for generic $c$ any trace on $\mathcal{O}_{c}$ is a restriction of a trace on $H_{c}$.
This implication of Theorem 0.4 (2) is proved in Corollary 6.4, in particular the Morita equivalence is given by bimodules $eH_{c}$ and $H_{c}e$. Now, for any two Morita equivalent algebras $A$ and $B$ with Morita equivalence given by an (AB)(A-B)-bimodule $M$ and a (BA)(B-A)-bimodule NN there is one-to-one correspondence between traces on $A$ and $B$ given by
$T_{B}\bigg(\sum_{i}n_{i}\otimes_{A}m_{i}\bigg)=T_{A}\bigg(\sum_{i}m_{i}\otimes_{B}n_{i}\bigg).$
The map $T_{B}$ is well-defined because $T_{A}$ is a trace and the map $T_{B}$ is a trace because $T_{A}$ is well-defined. Applying this to $A=H_{c}$, $B=\mathcal{O}_{c}$, we get
$T_{B}\bigg(\sum n_{i}\otimes_{A}m_{i}\bigg)=T_{A}\bigg(\sum m_{i}\otimes_{B}n_{i}\bigg)=T_{A}\bigg(\sum m_{i}n_{i}\bigg)=T_{A}\bigg(\sum n_{i}m_{i}\bigg),$
where we used that $T_{A}$ is a trace in the last step. Now, $\sum n_{i}\otimes_{A}m_{i}$ corresponds to just $\sum n_{i}m_{i}$, hence $T_{B}$ is the restriction of $T_{A}$ to $\mathcal{O}_{c}$.
For a classification of traces on $\mathcal{O}_{c}$, it therefore suffices to characterize traces on $H_{c}$.
@@
$gx=\varepsilon xg;\qquad gy=\varepsilon^{-1}yg;\qquad hx=yh;$
$hy=-xh;\qquad xy-yx=c=\sum_{i=0}^{n-1}c_{i}g^{i},\qquad\text{with $c_{0}\neq 0$},\ c_{i}=c_{-i}\text{ for all $i$}.$
Here the equality $c_{i}=c_{-i}$ holds because $c$ is in $Z(\mathbb{C}[\mathbb{D}_{n}])$ and the conjugacy classes of $\mathbb{D}_{n}$ are of the form $\big\{g^{i},g^{-i}\big\}$, $\big\{g^{2i}h\mid i\in\mathbb{Z}\big\}$, and $\big\{g^{2i+1}h\mid i\in\mathbb{Z}\big\}$.
We close this section with a definition of positive traces.
Definition 2.2.
For an algebra $B$ and an antilinear automorphism $\rho\colon B\to B$ such that $\rho^{2}=\operatorname{id}$, a trace $T$ on $B$ is said to be positive if $T(a\rho(a))>0$ for all nonzero $a$ in $B$.
The antilinear automorphism we study is defined on $H_{c}$ by $\rho(x)=y$, $\rho(y)=-x$, $\rho(g)=g$, $\rho(h)=h$. We return to positive traces in Section 6.
3 Algebra identities and relations
In this section, we introduce several important algebra elements and describe relations between them. The proofs are routine, but we provide them for the sake of completion.
Recall that $c=\sum_{i=0}^{n-1}c_{i}g^{i}$, with $g$ a generator of $\mathbb{D}_{n}$ such that $g^{n}$ is the group identity, and $\varepsilon=\exp\bigl(\frac{2\pi{\rm i}}{n}\bigr)$}. Let
@@ 3 Algebra identities and relations
$hk={}$ $\frac{1}{c_{0}}\left(hxy+\sum\limits_{i=1}^{n-1}\frac{hc_{i}g^{i}}{\varepsilon^{i}-1}\right)-\frac{h}{2}=\frac{1}{c_{0}}\left(-yxh+\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{-i}h}{\varepsilon^{i}-1}\right)-\frac{h}{2}$
$={}$ $\frac{1}{c_{0}}\left(c-xy+\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}}{\varepsilon^{-i}-1}\right)h-\frac{h}{2}=-\frac{1}{c_{0}}\left(xy+\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}}{\varepsilon^{i}-1}\right)h+\frac{h}{2}=-kh.$
@@
$[k,x]={}$ $\frac{1}{c_{0}}\left(xyx-x^{2}y+\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}x-xc_{i}g^{i}}{\varepsilon^{i}-1}\right)=\frac{1}{c_{0}}\left(-xc+\sum\limits_{i=1}^{n-1}\frac{\varepsilon^{i}xc_{i}g^{i}-xc_{i}g^{i}}{\varepsilon^{i}-1}\right)$
$={}$ $\frac{1}{c_{0}}\left(-xc+x\sum\limits_{i=1}^{n-1}c_{i}g^{i}\right)=-x.$
It follows immediately that kx=x(k1)kx=x(k-1), and the commutation relations with $y$ follow by conjugating the commutation relations with $x$ by $h$. ∎
@@
Clearly, $g^{i}$ and $e_{q}$ commute. We have
@@
$he_{q}=\frac{1}{n}\sum\limits_{i=0}^{n-1}\varepsilon^{iq}hg^{i}=\frac{1}{n}\sum\limits_{i=0}^{n-1}\varepsilon^{iq}g^{-i}h=e_{-q}h,$
@@
$e_{q}x=\frac{1}{n}\sum\limits_{i=0}^{n-1}\varepsilon^{iq}g^{i}x=\frac{1}{n}\sum\limits_{i=0}^{n-1}\varepsilon^{iq+i}xg^{i}=xe_{q+1},$
@@
$\alpha_{q}={}$ $\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{-iq}}{\varepsilon^{i}-1}=\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{-i}\varepsilon^{-iq}}{\varepsilon^{i}-1}=\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{iq}}{\varepsilon^{-i}-1}=\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{i(q+1)}}{1-\varepsilon^{i}}$
$={}$ $-\alpha_{-q-1}.$
@@
$\alpha_{p}e_{q}={}$ $\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{-ip}e_{q}}{\varepsilon^{i}-1}=\frac{1}{c_{0}}\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{i(q-p)}\varepsilon^{-iq}e_{q}}{\varepsilon^{i}-1}=\frac{1}{c_{0}}\left(\sum\limits_{i=1}^{n-1}\frac{c_{i}\varepsilon^{i(q-p)}g^{i}}{\varepsilon^{i}-1}\right)e_{q},$
@@
$k=\frac{1}{c_{0}}\Bigg(xy-\sum\limits_{q=0}^{n-1}\beta_{q}e_{q}\Bigg),\qquad\textrm{or}\qquad xy=c_{0}k+\sum\limits_{q=0}^{n-1}\beta_{q}e_{q},$
@@
$x^{a}y^{a}={}$ $\prod_{i=0}^{a-1}\Bigg(c_{0}(k+i)+\sum\limits_{q=0}^{n-1}\beta_{q+i}e_{q}\Bigg),$
$y^{a}x^{a}={}$ $(-1)^{a}\prod\limits_{i=0}^{a-1}\Bigg(c_{0}(-k+i)+\sum\limits_{q=0}^{n-1}\beta_{q+i}e_{-q}\Bigg).$
Also, the elements $x^{a}y^{a}e_{q}$ and $y^{a}x^{a}e_{q}$ can be written as polynomial expressions in $k$ multiplied by $e_{q}$ as follows
@@
We prove the polynomial expression for $x^{a}y^{a}$ by induction, with the base case of $a=1$ true by definition. Assume that $x^{a}y^{a}e_{q}=e_{q}\prod_{i=0}^{a-1}\bigl(c_{0}(k+i)+\sum_{q=0}^{n-1}\beta_{q+i}e_{q}\bigr)$. Then
$x^{a+1}y^{a+1}={}$ $x^{a}\Bigg(c_{0}k+\sum\limits_{q=0}^{n-1}\beta_{q}e_{q}\Bigg)y^{a}=x^{a}y^{a}\Bigg(c_{0}(k+a)+\sum\limits_{q=0}^{n-1}\beta_{q}e_{q-a}\Bigg)$
$={}$ $x^{a}y^{a}\Bigg(c_{0}(k+a)+\sum\limits_{q=0}^{n-1}\beta_{q+a}e_{q}\Bigg)=\prod\limits_{i=0}^{a}\Bigg(c_{0}(k+i)+\sum\limits_{q=0}^{n-1}\beta_{q+i}e_{q}\Bigg),$
as needed. The analogous claim for $y^{a}x^{a}$ follows from conjugation by $h$. The expressions for $x^{a}y^{a}e_{q}$ and $y^{a}x^{a}e_{q}$ follow using the statement that $e_{p}e_{q}=\delta_{p,q}e_{q}$ from Lemma 3.2. ∎
@@
By Lemma 3.3, we have
$xye_{q}={}$ $\Bigg(c_{0}k+\frac{c_{0}}{2}-\sum\limits_{i=1}^{n-1}\frac{c_{i}g^{i}}{\varepsilon^{i}-1}\Bigg)e_{q}=\left(c_{0}k+\frac{c_{0}}{2}-c_{0}\alpha_{q}\right)e_{q}.$
Also, by definition, $xye_{q}=(c_{0}k+\beta_{q})e_{q}$. Thus
$\left(c_{0}k+\beta_{q}\right)e_{q}=\left(c_{0}k+\frac{c_{0}}{2}-c_{0}\alpha_{q}\right)e_{q},\qquad\beta_{q}e_{q}=\left(\frac{c_{0}}{2}-c_{0}\alpha_{q}\right)e_{q},$
@@
4 Characterization of traces
In this section, we study traces using basis elements of the form $x^{i}R(k)e_{q}$, $y^{i}R(k)e_{q}$, $x^{i}R(k)e_{q}h$, and $y^{i}R(k)e_{q}h$, with $R(k)\in\mathbb{C}[k]$. Let $T$ be a trace on $H_{c}$. We begin with a lemma about the values $T(e_{q}h)$.
Lemma 4.1.
If $m\nmid q$, then $T(e_{q}h)=0$.